Paragraph sded with the Interature with having the isomer of the m e fact of the largest generated alkene for an E1 reaction is the one possessing the highest degree of substitution. Thus it is the most stable product with the lowest energy most of the product isomers. By obtaining these results it was determined that the overall experiment was indeed and a success. Questions- : he ane 27, 6-56. Gas chromatographic analysis of a mixture of organic compounds gave the following peak areas cm 1.6, (a) Calculate the mole percent composition of the mixture. Assume that the response of the as in
Answer #1

Gas chromatography

(a) mole percent

Total area = 6.6

mole percent hexane = 2.7 x 100/6.6 = 0.41%

mole percent heptane = 1.6 x 100/6.6 = 0.24%

mole percent hexanol = 1.8 x 100/6.6 = 0.27%

mole percent toluene = 0.5 x 100/6.6 = 0.08%

(b) mass percent

mass hexane = 2.7 x 86.18 g/mol = 233 g

mass heptane = 1.6 x 100.21 g/mol = 160.3 g

mass hexanol = 1.8 x 102.2 g/mol = 184 g

mass toluene = 0.5 x 92.14 g/mol = 46.07 g

Total mass = 623.4 g

mass percent hexane = 233 x 100/623.4 = 37.4%

mass percent heptane = 160.3 x 100/623.4 = 25.7%

mass percent hexanol = 184 x 100/623.4 = 29.5%

mass percent toluene = 46.07 x 100/623.4 = 7.4%